investment is approximately equal to 1.11 The average cost of a WV public four-year, full-time, in-state Colleges tuition and fees was $4,548.44 in 2010, and is $7,056,00 in 2019. The annual rate of increase is approximately
Answer #1

F = P(1 + i)n

7056 = 4548.44(1 + i)9

(1 + i)9 = 7056 / 4548.44 = 1.5513

1 + i = 9th root of 1.5513

1 + i = 1.05

i = 1 - 1.05 = 0.05 or 5%

Thus, the annual rate of increase is approximately 5%.